OpenStudy (anonymous):

multiply and simplify: sqrt 2(sqrt 8+4 sqrt 3)

OpenStudy (anonymous):

damn equation button is broken again

OpenStudy (ash2326):

Is this your question \[\sqrt 2(\sqrt8+4\sqrt3) \]?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

I assume it is let's use the distributive property of multiplication A(B+C)=AB+AC we get \[\sqrt {16}+4\sqrt{6}\] we know 16=4*4 so we get \[4+4\sqrt6\] finally we get \[4(1+\sqrt 6)\]